LBT/MODS1

PI: Nathan Smith

Observer: Wen-fai Fong

03:25 UT: Start test science script.

03:30 UT: Start observing 1st target for imaging, ASASSN-15ua. airmass=1.2, seeing=0.84".

03:57 UT: Start spectroscopy script, ASASSN-15ua. airmass=1.3, seeing=1.0".

04:30 UT: Start spectroscopy for 2nd target, NGC5775OT. airmass=1.2, seeing=1.0".

06:10 UT: Slew to standard star.

06:16 UT: Start spectroscopy script for standard star BD+33D2642. airmass=1.1, seeing=1.1". After taking and reading out 3 exposures (checked this in /newdata), script hung "Red Exposures Done, Blue Channel still acquiring." Tried typing "red expdone blue expdone" in the GUI but it didn't do anything. So just aborted script.

06:45 UT: Start spectroscopy for 3rd target, ASASSN-16ex. airmass=1.1, seeing=0.91".

07:00 UT: Start imaging of 4th target, SN2010mc. airmass=1.1, seeing=0.89".

07:30 UT: Start spectroscopy of 4th target, SN2010mc. airmass=1.1, seeing=1.25".

08:45 UT: Red channel exposure has strange pattern (mods1r....0037.fits). After image 37, red channel images no longer appear in newdata.

09:20 UT: After 6 exposures in a row not reading out in the red channel, paused script until further notice.

Could not trouble shoot the problem. Would require full restart of MODS which would take us into sunrise. Lost the rest of the night.

-- %USERSIG{EiichiEgami - 2016-07-09}%

Comments


 

-- %USERSIG{EiichiEgami - 2016-07-09}%

Comments

 
Topic revision: r3 - 09 Jul 2016, EiichiEgami
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback